# Flexible and efficient spatial extremes emulation via variational autoencoders

### v3

@@ B.1 Properties of exponentially-tilted positive-stable variables
$\begin{split}\chi_{ij}&=\lim_{u\rightarrow 1}\frac{\Pr\{X(\boldsymbol{s}_{i})>F_{i}^{-1}(u),X(\boldsymbol{s}_{j})>F_{j}^{-1}(u)\}}{1-u}\\ &=\lim_{t\rightarrow\infty}t\Pr\{X(\boldsymbol{s}_{i})>q_{i}(t),X(\boldsymbol{s}_{j})>q_{j}(t)\}\\ &=\lim_{t\rightarrow\infty}t\left[1-2\left(1-\frac{1}{t}\right)+\Pr\{X(\boldsymbol{s}_{i})\leq q_{i}(t),X(\boldsymbol{s}_{j})\leq q_{j}(t))\}\right]\\ &=\lim_{t\rightarrow\infty}2-t\left[1-F_{ij}\{q_{i}(t),q_{j}(t)\}\right],\end{split}$ (B.8)
@@
$\begin{split}\lim_{t\rightarrow\infty}\frac{\log\Pr\{X(\boldsymbol{s}_{i})>q_{i}(t),X(\boldsymbol{s}_{j})>q_{j}(t)\}}{\log t}=-\frac{1}{\eta_{ij}},\end{split}$ (B.9)
@@
    If $\mathcal{C}_{i}\cap\mathcal{D}=\emptyset$ and $\mathcal{C}_{j}\cap\mathcal{D}=\emptyset$, we know from Corollary 1.1.1 that $q_{i}(t)=c_{i}^{\alpha_{0}}t^{\alpha_{0}}\{1+R_{i}(t)+o(t^{-1})\}$ and $q_{j}(t)=c_{j}^{\alpha_{0}}t^{\alpha_{0}}\{1+R_{j}(t)+o(t^{-1})\}$, in which $R_{i}(t)=\alpha_{0}({d_{i}}/{c_{i}^{2}}-1/2)t^{-1}$ and $R_{j}(t)=\alpha_{0}({d_{j}}/{c_{j}^{2}}-1/2)t^{-1}$. Similarly to Proposition 1.1, we first deduce
    $\log F_{ij}$$\{q_{i}(t),q_{j}(t)\}=\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha}-\sum_{k=1}^{K}\left[\gamma_{k}+\frac{\tau^{1/\alpha_{0}}\omega_{ki}^{1/\alpha}}{c_{i}t\{1+R_{i}(t)+o(t^{-1})\}}+\frac{\tau^{1/\alpha_{0}}\omega_{kj}^{1/\alpha}}{c_{j}t\{1+R_{j}(t)+o(t^{-1})\}}\right]^{\alpha}$
    $=\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha}-\sum_{k\in\bar{\mathcal{D}}}\left[\gamma_{k}+\frac{\tau^{1/\alpha_{0}}\omega_{ki}^{1/\alpha}}{c_{i}t}\{1-R_{i}(t)\}+\frac{\tau^{1/\alpha_{0}}\omega_{kj}^{1/\alpha}}{c_{j}t}\{1-R_{j}(t)\}+o\left(\frac{1}{t^{2}}\right)\right]^{\alpha}$
    $=\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha}-\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha}\left[1+\frac{\alpha\tau^{1/\alpha_{0}}\omega_{ki}^{1/\alpha}/\gamma_{k}}{c_{i}t}\{1-R_{i}(t)\}+\frac{\alpha\tau^{1/\alpha_{0}}\omega_{kj}^{1/\alpha}/\gamma_{k}}{c_{j}t}\{1-R_{j}(t)\}+o\left(\frac{1}{t^{2}}\right)\right],$
    in which the Taylor expansion in Eq. (B.2) is applied. Recall the definitions of $c_{i}$ and $c_{j}$ in Proposition 1.1, and we find
@@
    Then it follows from Eq. (B.4) that
@@
    Plugging this result into (B.8), we have $\chi_{ij}=\lim_{t\rightarrow\infty}R_{i}(t)+R_{j}(t)+o(t^{-1})=0$.
@@
    $\log\Pr\{X(\boldsymbol{s}_{i})>q_{i}(t),X(\boldsymbol{s}_{j})>q_{j}(t)\}\sim\log\frac{R_{i}(t)+R_{j}(t)}{t}=\log\alpha_{0}\left(\frac{d_{i}}{c_{i}^{2}}+\frac{d_{j}}{c_{j}^{2}}-1\right)-2\log t$
    as $t\rightarrow\infty$. By Eq. (B.9), $\eta_{ij}=1/2$.
@@
    If $\mathcal{C}_{i}\cap\mathcal{D}=\emptyset$ and $\mathcal{C}_{j}\cap\mathcal{D}\neq\emptyset$, we know from Corollary 1.1.1 that $q_{i}(t)\sim c_{i}{}^{\alpha_{0}}t^{\alpha_{0}}\{1+R_{i}(t)+o(t^{-1})\}$ and $q_{j}(t)\sim c^{\prime}_{j}{}^{\alpha_{0}/\alpha}t^{\alpha_{0}/\alpha}\{1+R^{*}_{j}(t)+O(t^{-1/\alpha})\}$ as $t\rightarrow\infty$, in which $R_{i}(t)=\alpha_{0}({d_{i}}/{c_{i}^{2}}-1/2)t^{-1}$ and $R^{*}_{j}(t)=\alpha_{0}c_{j}t^{1-1/\alpha}/(\alpha c^{\prime}_{j}{}^{1/\alpha})-\alpha_{0}t^{-1}/(2\alpha)$. Then
@@
    $\begin{split}\log&F_{ij}\{q_{i}(t),q_{j}(t)\}=\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha}-\sum_{k=1}^{K}\left\{\gamma_{k}+\frac{\tau^{1/\alpha_{0}}\omega_{ki}^{1/\alpha}}{q_{i}^{1/\alpha_{0}}(t)}+\frac{\tau^{1/\alpha_{0}}\omega_{kj}^{1/\alpha}}{q_{j}^{1/\alpha_{0}}(t)}\right\}^{\alpha}\\ =&\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha}-\sum_{k\in\bar{\mathcal{D}}}\left\{\gamma_{k}+\frac{\tau^{1/\alpha_{0}}\omega_{ki}^{1/\alpha}}{q_{i}^{1/\alpha_{0}}(t)}+\frac{\tau^{1/\alpha_{0}}\omega_{kj}^{1/\alpha}}{q_{j}^{1/\alpha_{0}}(t)}\right\}^{\alpha}-\sum_{k\in\mathcal{D}}\frac{\tau^{\alpha/\alpha_{0}}\omega_{kj}}{q_{j}^{\alpha/\alpha_{0}}(t)}\\ =&\log F_{j}\{q_{j}(t)\}+\sum_{k\in\bar{\mathcal{D}}}\left\{\gamma_{k}+\frac{\tau^{1/\alpha_{0}}\omega_{kj}^{1/\alpha}}{q_{j}^{1/\alpha_{0}}(t)}\right\}^{\alpha}-\sum_{k\in\bar{\mathcal{D}}}\left\{\gamma_{k}+\frac{\tau^{1/\alpha_{0}}\omega_{ki}^{1/\alpha}}{q_{i}^{1/\alpha_{0}}(t)}+\frac{\tau^{1/\alpha_{0}}\omega_{kj}^{1/\alpha}}{q_{j}^{1/\alpha_{0}}(t)}\right\}^{\alpha}.\end{split}$ (B.10)
    For the two summations on the right, we split $\bar{\mathcal{D}}$ into $\mathcal{C}_{j}\cap\bar{\mathcal{D}}$ and $\bar{\mathcal{D}}\setminus(\mathcal{C}_{j}\cap\bar{\mathcal{D}})$:
    $\begin{split}&\sum_{k\in\mathcal{C}_{j}\cap\bar{\mathcal{D}}}\left\{\gamma_{k}+\frac{\tau^{1/\alpha_{0}}\omega_{kj}^{1/\alpha}}{q_{j}^{1/\alpha_{0}}(t)}\right\}^{\alpha}-\sum_{k\in\mathcal{C}_{j}\cap\bar{\mathcal{D}}}\left\{\gamma_{k}+\frac{\tau^{1/\alpha_{0}}\omega_{ki}^{1/\alpha}}{q_{i}^{1/\alpha_{0}}(t)}+\frac{\tau^{1/\alpha_{0}}\omega_{kj}^{1/\alpha}}{q_{j}^{1/\alpha_{0}}(t)}\right\}^{\alpha}=\\ &-\sum_{k\in\mathcal{C}_{j}\cap\bar{\mathcal{D}}}\frac{\alpha\tau^{1/\alpha_{0}}\omega_{ki}^{1/\alpha}}{\gamma_{k}^{1-\alpha}q_{i}^{1/\alpha_{0}}(t)}-\frac{\alpha(\alpha-1)}{2}\sum_{k\in\mathcal{C}_{j}\cap\bar{\mathcal{D}}}\frac{\tau^{2/\alpha_{0}}}{\gamma_{k}^{2-\alpha}}\left\{\frac{\omega_{ki}^{2/\alpha}}{q_{i}^{2/\alpha_{0}}(t)}+\frac{2\omega_{ki}^{1/\alpha}\omega_{kj}^{1/\alpha}}{q_{i}^{1/\alpha_{0}}(t)q_{j}^{1/\alpha_{0}}(t)}\right\}+o\left(t^{-1-\frac{1}{\alpha}}\right)\end{split}$ (B.11)
    and
    $\begin{split}\sum_{k\in\bar{\mathcal{D}}\setminus(\mathcal{C}_{j}\cap\bar{\mathcal{D}})}&\left\{\gamma_{k}+\frac{\tau^{1/\alpha_{0}}\omega_{kj}^{1/\alpha}}{q_{j}^{1/\alpha_{0}}(t)}\right\}^{\alpha}-\sum_{k\in\bar{\mathcal{D}}\setminus(\mathcal{C}_{j}\cap\bar{\mathcal{D}})}\left\{\gamma_{k}+\frac{\tau^{1/\alpha_{0}}\omega_{ki}^{1/\alpha}}{q_{i}^{1/\alpha_{0}}(t)}+\frac{\tau^{1/\alpha_{0}}\omega_{kj}^{1/\alpha}}{q_{j}^{1/\alpha_{0}}(t)}\right\}^{\alpha}\\ &\sum_{k\in\bar{\mathcal{D}}\setminus(\mathcal{C}_{j}\cap\bar{\mathcal{D}})}\gamma_{k}^{\alpha}-\sum_{k\in\bar{\mathcal{D}}\setminus(\mathcal{C}_{j}\cap\bar{\mathcal{D}})}\left\{\gamma_{k}+\frac{\tau^{1/\alpha_{0}}\omega_{ki}^{1/\alpha}}{q_{i}^{1/\alpha_{0}}(t)}\right\}^{\alpha}.\\ \end{split}$ (B.12)
    Feeding Eqs. (B.11) and (B.12) back into (B.10), we have
@@
    $\begin{split}\log F_{ij}\{q_{i}(t),q_{j}(t)\}&=\log F_{j}\{q_{j}(t)\}+\log F_{i}\{q_{i}(t)\}+\frac{c_{ij}}{q_{i}^{1/\alpha_{0}}(t)q_{j}^{1/\alpha_{0}}(t)}+o\left(t^{-1-\frac{1}{\alpha}}\right)\text{ as }t\rightarrow\infty,\end{split}$
    in which $c_{ij}=\alpha(\alpha-1)\tau^{2/\alpha_{0}}\sum_{k\in\mathcal{C}_{i}\cap\mathcal{C}_{j}}\gamma_{k}^{\alpha-2}\omega_{ki}^{1/\alpha}\omega_{kj}^{1/\alpha}$.
    Then it follows from Eqs. (B.4) and (B.8) that
    $1-F_{ij}\{q_{i}(t),q_{j}(t)\}=$ $1-\exp\left[\log F_{j}\{q_{j}(t)\}+\log F_{i}\{q_{i}(t)\}+\frac{c_{ij}}{q_{i}^{1/\alpha_{0}}(t)q_{j}^{1/\alpha_{0}}(t)}+o\left(t^{-1-\frac{1}{\alpha}}\right)\right]$
    $=$ $-\log F_{j}\{q_{j}(t)\}-\log F_{i}\{q_{i}(t)\}-\frac{c_{ij}}{q_{i}^{1/\alpha_{0}}(t)q_{j}^{1/\alpha_{0}}(t)}+o\left(t^{-1-\frac{1}{\alpha}}\right)$
    $=$ $-\log\left(1-\frac{1}{t}\right)-\log\left(1-\frac{1}{t}\right)-\frac{c_{ij}}{q_{i}^{1/\alpha_{0}}(t)q_{j}^{1/\alpha_{0}}(t)}+o\left(t^{-1-\frac{1}{\alpha}}\right)$
@@
    $=$ $\frac{2}{t}-\frac{c_{ij}}{c_{i}c^{\prime}_{j}{}^{1/\alpha}t^{1+1/\alpha}}+o\left(t^{-1-\frac{1}{\alpha}}\right)$
    Therefore, $\Pr\{X(\boldsymbol{s}_{i})>q_{i}(t),X(\boldsymbol{s}_{j})>q_{j}(t)\}=\frac{c_{ij}}{c_{i}c^{\prime}_{j}{}^{1/\alpha}}t^{-1-1/\alpha}+o\left(t^{-1-1/\alpha}\right)$. Then it follows from Eqs. (B.8) and (B.9) that $\xi_{ij}=0$ and $\eta_{ij}=1/2$.
@@
    When $\mathcal{C}_{i}\cap\mathcal{D}\neq\emptyset$ and $\mathcal{C}_{j}\cap\mathcal{D}\neq\emptyset$, we have $q_{i}(t)\sim c^{\prime}_{i}{}^{\alpha_{0}/\alpha}t^{\alpha_{0}/\alpha}\left\{1+O(t^{1-1/\alpha})\right\}$ and $q_{j}(t)\sim c^{\prime}_{j}{}^{\alpha_{0}/\alpha}t^{\alpha_{0}/\alpha}\left\{1+O(t^{1-1/\alpha})\right\}$ as $t\rightarrow\infty$. Then

### v4

@@ A.1 Properties of exponentially-tilted positive-stable variables
$\begin{split}\chi_{ij}&=\lim_{u\rightarrow 1}\frac{\Pr\{X(\boldsymbol{s}_{i})>F_{i}^{-1}(u),X(\boldsymbol{s}_{j})>F_{j}^{-1}(u)\}}{1-u}\\ &=\lim_{t\rightarrow\infty}t\Pr\{X(\boldsymbol{s}_{i})>q_{i}(t),X(\boldsymbol{s}_{j})>q_{j}(t)\}\\ &=\lim_{t\rightarrow\infty}t\left[1-2\left(1-\frac{1}{t}\right)+\Pr\{X(\boldsymbol{s}_{i})\leq q_{i}(t),X(\boldsymbol{s}_{j})\leq q_{j}(t))\}\right]\\ &=\lim_{t\rightarrow\infty}2-t\left[1-F_{ij}\{q_{i}(t),q_{j}(t)\}\right],\end{split}$ (A.8)
@@
$\begin{split}\lim_{t\rightarrow\infty}\frac{\log\Pr\{X(\boldsymbol{s}_{i})>q_{i}(t),X(\boldsymbol{s}_{j})>q_{j}(t)\}}{\log t}=-\frac{1}{\eta_{ij}},\end{split}$ (A.9)
@@
To facilitate the proofs of each case listed in Theorem 3.3, we first introduce some constants for simplicity:
$\begin{split}c_{ij}&=\alpha(\alpha-1)\sum_{k\in\mathcal{C}_{i}\cap\mathcal{C}_{j}}\gamma_{k}^{\alpha-2}\omega_{ki}^{1/\alpha}\omega_{kj}^{1/\alpha},\text{ and }d_{ij}=\sum_{k\in\mathcal{D}}\left(\frac{\omega_{ki}^{1/\alpha}}{c^{\prime}_{i}{}^{1/\alpha}}+\frac{\omega_{kj}^{1/\alpha}}{c^{\prime}_{j}{}^{1/\alpha}}\right)^{\alpha}.\end{split}$ (A.10)
In addition, constants $c_{j}$, $c_{j}^{\prime}$ and $d_{j}$ are defined in Eq. (A.2).
@@
    If $\mathcal{C}_{i}\cap\mathcal{D}=\emptyset$ and $\mathcal{C}_{j}\cap\mathcal{D}=\emptyset$, we know from Corollary A.1.1 that
    $\begin{split}q_{i}(t)&=\tau c_{i}^{\alpha_{0}}t^{\alpha_{0}}\{1+R_{i}(t)+o(t^{-1})\},\\ q_{j}(t)&=\tau c_{j}^{\alpha_{0}}t^{\alpha_{0}}\{1+R_{j}(t)+o(t^{-1})\},\end{split}$
    in which $R_{i}(t)=\alpha_{0}({d_{i}}/{c_{i}^{2}}-1/2)t^{-1}$ and $R_{j}(t)=\alpha_{0}({d_{j}}/{c_{j}^{2}}-1/2)t^{-1}$. Using the joint distribution in Proposition 3.2 in the main paper, we first deduce
    $\log F_{ij}$$\{q_{i}(t),q_{j}(t)\}=\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha}-\sum_{k=1}^{K}\left[\gamma_{k}+\frac{\omega_{ki}^{1/\alpha}}{c_{i}t\{1+R_{i}(t)+o(t^{-1})\}}+\frac{\omega_{kj}^{1/\alpha}}{c_{j}t\{1+R_{j}(t)+o(t^{-1})\}}\right]^{\alpha}$
    $=\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha}-\sum_{k\in\bar{\mathcal{D}}}\left[\gamma_{k}+\frac{\omega_{ki}^{1/\alpha}}{c_{i}t}\{1-R_{i}(t)\}+\frac{\omega_{kj}^{1/\alpha}}{c_{j}t}\{1-R_{j}(t)\}+o\left(\frac{1}{t^{2}}\right)\right]^{\alpha}$
    $=\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha}-\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha}\left[1+\frac{\alpha\omega_{ki}^{1/\alpha}/\gamma_{k}}{c_{i}t}\{1-R_{i}(t)\}+\frac{\alpha\omega_{kj}^{1/\alpha}/\gamma_{k}}{c_{j}t}\{1-R_{j}(t)\}+o\left(\frac{1}{t^{2}}\right)\right],$
    in which the penultimate equality uses the negative binomial expansion and the last euqaltiy is derived from the Taylor expansion in Eq. (A.3). Recall the definitions of $c_{i}$ and $c_{j}$ in Proposition 3.1 from the main paper, and we find
@@
    Then it follows from Eq. (A.4) that
@@
    Plugging this result into (A.8), we have $\chi_{ij}=\lim_{t\rightarrow\infty}\{R_{i}(t)+R_{j}(t)+o(t^{-1})\}=0$.
@@
    $\log\Pr\{X(\boldsymbol{s}_{i})>q_{i}(t),X(\boldsymbol{s}_{j})>q_{j}(t)\}$ $\sim\log\frac{R_{i}(t)+R_{j}(t)}{t}$
    $=\log\alpha_{0}+\log\left(\frac{d_{i}}{c_{i}^{2}}+\frac{d_{j}}{c_{j}^{2}}-1\right)-2\log t$
    as $t\rightarrow\infty$. By Eq. (A.9), $\eta_{ij}=1/2$.
@@
    If $\mathcal{C}_{i}\cap\mathcal{D}=\emptyset$ and $\mathcal{C}_{j}\cap\mathcal{D}\neq\emptyset$, we deduce $c_{i}\neq 0$, $c_{i}^{\prime}=0$, $c_{j}^{\prime}\neq 0$ and $\mathcal{C}_{i}=\bar{\mathcal{D}}$. From Corollary A.1.1,
    $\begin{split}q_{i}(t)&\sim\tau c_{i}{}^{\alpha_{0}}t^{\alpha_{0}}\{1+R_{i}(t)+o(t^{-1})\},\\ q_{j}(t)&\sim\tau c^{\prime}_{j}{}^{\alpha_{0}/\alpha}t^{\alpha_{0}/\alpha}\{1+R^{*}_{j}(t)+O(t^{-1/\alpha})\},\end{split}$ (A.11)
    as $t\rightarrow\infty$, where $R_{i}(t)=\alpha_{0}({d_{i}}/{c_{i}^{2}}-1/2)t^{-1}$ and $R^{*}_{j}(t)=\alpha_{0}c_{j}t^{1-1/\alpha}/(\alpha c^{\prime}_{j}{}^{1/\alpha})-\alpha_{0}t^{-1}/(2\alpha)$. Again by the joint distribution in Proposition 3.2 of the main paper,
@@
    $\begin{split}\log&F_{ij}\{q_{i}(t),q_{j}(t)\}=\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha}-\sum_{k=1}^{K}\left\{\gamma_{k}+\frac{\tau^{1/\alpha_{0}}\omega_{ki}^{1/\alpha}}{q_{i}^{1/\alpha_{0}}(t)}+\frac{\tau^{1/\alpha_{0}}\omega_{kj}^{1/\alpha}}{q_{j}^{1/\alpha_{0}}(t)}\right\}^{\alpha}\\ =&\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha}-\sum_{k\in\bar{\mathcal{D}}}\left\{\gamma_{k}+\frac{\tau^{1/\alpha_{0}}\omega_{ki}^{1/\alpha}}{q_{i}^{1/\alpha_{0}}(t)}+\frac{\tau^{1/\alpha_{0}}\omega_{kj}^{1/\alpha}}{q_{j}^{1/\alpha_{0}}(t)}\right\}^{\alpha}-\sum_{k\in\mathcal{D}}\frac{\tau^{\alpha/\alpha_{0}}\omega_{kj}}{q_{j}^{\alpha/\alpha_{0}}(t)}.\end{split}$
    Here, we split the sum over $k\in{1,\ldots,K}$ to $k\in\mathcal{D}$ and $k\in\bar{\mathcal{D}}$. The third summation can be re-written as $\sum_{k\in\mathcal{D}}\frac{\tau^{\alpha/\alpha_{0}}\omega_{kj}}{q_{j}^{\alpha/\alpha_{0}}(t)}=c_{j}^{\prime}\left\{\frac{q_{j}(t)}{\tau}\right\}^{-\alpha/\alpha_{0}}$. For the second summation, we apply Eq. (A.3) again to get
    $\sum_{k\in\bar{\mathcal{D}}}$$\left\{\gamma_{k}+\frac{\tau^{1/\alpha_{0}}\omega_{ki}^{1/\alpha}}{q_{i}^{1/\alpha_{0}}(t)}+\frac{\tau^{1/\alpha_{0}}\omega_{kj}^{1/\alpha}}{q_{j}^{1/\alpha_{0}}(t)}\right\}^{\alpha}$
    $=$ $\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha}\left[1+\frac{\alpha\tau^{1/\alpha_{0}}\omega_{ki}^{1/\alpha}}{\gamma_{k}q_{i}^{1/\alpha_{0}}(t)}+\frac{\alpha\tau^{1/\alpha_{0}}\omega_{kj}^{1/\alpha}}{\gamma_{k}q_{j}^{1/\alpha_{0}}(t)}+\frac{\alpha(\alpha-1)}{2\gamma_{k}^{2}}\left\{\frac{\tau^{1/\alpha_{0}}\omega_{ki}^{1/\alpha}}{q_{i}^{1/\alpha_{0}}(t)}+\frac{\tau^{1/\alpha_{0}}\omega_{kj}^{1/\alpha}}{q_{j}^{1/\alpha_{0}}(t)}\right\}^{2}+o(t^{-1-\frac{1}{\alpha}})\right]$
    $=$ $\sum_{k\in\bar{\mathcal{D}}}\gamma_{k}^{\alpha}+c_{i}\left\{\frac{q_{i}(t)}{\tau}\right\}^{-\frac{1}{\alpha_{0}}}+c_{j}\left\{\frac{q_{j}(t)}{\tau}\right\}^{-\frac{1}{\alpha_{0}}}+d_{i}\left\{\frac{q_{i}(t)}{\tau}\right\}^{-\frac{2}{\alpha_{0}}}+d_{j}\left\{\frac{q_{j}(t)}{\tau}\right\}^{-\frac{2}{\alpha_{0}}}$
    $+c_{ij}\left\{\frac{q_{i}(t)q_{j}(t)}{\tau^{2}}\right\}^{-\frac{1}{\alpha_{0}}}+o(t^{-1-\frac{1}{\alpha}}),$ (A.12)
    in which the constants $c_{i}$, $c_{j}$, $d_{i}$, $d_{j}$ and $c_{ij}$ are defined previously in Eq. (A.10), and the residual term $o\left(t^{-\frac{2}{\alpha}}\right)$ is derived using the asymptotic approximation of $q_{i}(t)$ and $q_{j}(t)$ in Eq. (A.11).
    Combining this result with Eq. (A.12) and feeding them into Eq. (A.14), we have
    11 11 $-F_{ij}\{q_{i}(t),q_{j}(t)\}=$
    $=1-\exp\left[c_{i}\left\{\frac{q_{i}(t)}{\tau}\right\}^{-\frac{1}{\alpha_{0}}}+c_{j}\left\{\frac{q_{j}(t)}{\tau}\right\}^{-\frac{1}{\alpha_{0}}}+c_{j}^{\prime}\left\{\frac{q_{j}(t)}{\tau}\right\}^{-\frac{\alpha}{\alpha_{0}}}\right.$
    $\left.+d_{i}\left\{\frac{q_{i}(t)}{\tau}\right\}^{-\frac{2}{\alpha_{0}}}+d_{j}\left\{\frac{q_{j}(t)}{\tau}\right\}^{-\frac{2}{\alpha_{0}}}+c_{ij}\left\{\frac{q_{i}(t)q_{j}(t)}{\tau^{2}}\right\}^{-\frac{1}{\alpha_{0}}}+o\left(t^{-\frac{2}{\alpha}}\right)\right]$
    $=$ $c_{i}\left\{\frac{q_{i}(t)}{\tau}\right\}^{-\frac{1}{\alpha_{0}}}+c_{j}\left\{\frac{q_{j}(t)}{\tau}\right\}^{-\frac{1}{\alpha_{0}}}+c_{j}^{\prime}\left\{\frac{q_{j}(t)}{\tau}\right\}^{-\frac{\alpha}{\alpha_{0}}}+d_{i}\left\{\frac{q_{i}(t)}{\tau}\right\}^{-\frac{2}{\alpha_{0}}}+d_{j}\left\{\frac{q_{j}(t)}{\tau}\right\}^{-\frac{2}{\alpha_{0}}}$
    $+c_{ij}\left\{\frac{q_{i}(t)q_{j}(t)}{\tau^{2}}\right\}^{-\frac{1}{\alpha_{0}}}-\frac{c_{i}^{2}}{2}\left\{\frac{q_{i}(t)}{\tau}\right\}^{-\frac{2}{\alpha_{0}}}-\frac{c_{j}^{2}}{2}\left\{\frac{q_{j}(t)}{\tau}\right\}^{-\frac{2}{\alpha_{0}}}-\frac{c_{j}^{\prime 2}}{2}\left\{\frac{q_{j}(t)}{\tau}\right\}^{-\frac{2\alpha}{\alpha_{0}}}$
    $-c_{i}c_{j}\left\{\frac{q_{i}(t)q_{j}(t)}{\tau^{2}}\right\}^{-\frac{1}{\alpha_{0}}}-c_{i}c_{j}^{\prime}\frac{q_{i}^{-\frac{1}{\alpha_{0}}}(t)q^{-\frac{\alpha}{\alpha_{0}}}_{j}(t)}{\tau^{-\frac{1+\alpha}{\alpha_{0}}}}-c_{j}c_{j}^{\prime}\frac{q_{j}^{-\frac{1+\alpha}{\alpha_{0}}}(t)}{\tau^{-\frac{1+\alpha}{\alpha_{0}}}}+o(t^{-1-\frac{1}{\alpha}}).$
    Then we utilize the asymptotic approximation of the marginal distribution in Eq. (A.2) to get
    $\begin{split}1-F_{ij}\{q_{i}(t),&q_{j}(t)\}=\bar{F}_{i}\{q_{i}(t)\}+\bar{F}_{j}\{q_{j}(t)\}+c_{ij}\left\{\frac{q_{i}(t)q_{j}(t)}{\tau^{2}}\right\}^{-\frac{1}{\alpha_{0}}}\\ &-c_{i}c_{j}\left\{\frac{q_{i}(t)q_{j}(t)}{\tau^{2}}\right\}^{-\frac{1}{\alpha_{0}}}-c_{i}c_{j}^{\prime}\frac{q_{i}^{-\frac{1}{\alpha_{0}}}(t)q^{-\frac{\alpha}{\alpha_{0}}}_{j}(t)}{\tau^{-\frac{1+\alpha}{\alpha_{0}}}}-c_{j}c_{j}^{\prime}\frac{q_{j}^{-\frac{1+\alpha}{\alpha_{0}}}(t)}{\tau^{-\frac{1+\alpha}{\alpha_{0}}}}+o(t^{-1-\frac{1}{\alpha}}).\end{split}$
    By definition, $t^{-1}=\bar{F}_{i}\{q_{i}(t)\}=\bar{F}_{j}\{q_{j}(t)\}$. Therefore,
    $\chi_{ij}=\lim_{t\rightarrow\infty}2-t\left[1-F_{ij}\{q_{i}(t),q_{j}(t)\}\right]=0.$
    If $\mathcal{C}_{i}\cap\mathcal{C}_{j}=\emptyset$, $c_{ij}=0$ and $c_{j}=0$. By Eq. (A.9) and (A.11),
@@
    $\begin{split}-\frac{1}{\eta_{ij}}=\lim_{t\rightarrow\infty}\frac{\log\left[c_{i}c_{j}^{\prime}\frac{q_{i}^{-{1}/{\alpha_{0}}}(t)q^{-{\alpha}/{\alpha_{0}}}_{j}(t)}{\tau^{-{1+\alpha}/{\alpha_{0}}}}+o\left(t^{-1-\frac{1}{\alpha}}\right)\right]}{\log t}=-2,\end{split}$
    and $\eta_{ij}=1/2$.
    If $\mathcal{C}_{i}\cap\mathcal{C}_{j}\neq\emptyset$, then $\bar{\mathcal{D}}\cap\mathcal{C}_{j}\neq\emptyset$ and $c_{j}>0$, $c_{ij}<0$. Thus, $02c_{i}c_{j}-c_{ij}>$ 0 and by Eq. (A.11),
    $c_{i}c_{j}\left\{\frac{q_{i}(t)q_{j}(t)}{\tau^{2}}\right\}^{-\frac{1}{\alpha_{0}}}+c_{j}c_{j}^{\prime}\frac{q_{j}^{-\frac{1+\alpha}{\alpha_{0}}}(t)}{\tau^{-\frac{1+\alpha}{\alpha_{0}}}}-c_{ij}\left\{\frac{q_{i}(t)q_{j}(t)}{\tau^{2}}\right\}^{-\frac{1}{\alpha_{0}}}=\frac{2c_{i}c_{j}-c_{ij}}{c_{i}c_{j}^{\prime}{}^{1/\alpha}}t^{-1-\frac{1}{\alpha}}+o(t^{-1-\frac{1}{\alpha}}).$
    Consequently,
@@
    $\begin{split}-\frac{1}{\eta_{ij}}=\lim_{t\rightarrow\infty}\frac{\log\left[\frac{2c_{i}c_{j}-c_{ij}}{c_{i}c_{j}^{\prime}{}^{1/\alpha}}t^{-1-\frac{1}{\alpha}}+o(t^{-1-\frac{1}{\alpha}})\right]}{\log t}=-1-\frac{1}{\alpha},\end{split}$
    and $\eta_{ij}=\alpha/(\alpha+1)$.
@@
    When $\mathcal{C}_{i}\cap\mathcal{D}\neq\emptyset$ and $\mathcal{C}_{j}\cap\mathcal{D}\neq\emptyset$, we have $c^{\prime}_{i}\neq 0$, $c^{\prime}_{j}\neq 0$ and
    $\begin{split}q_{i}(t)&\sim\tau c^{\prime}_{i}{}^{\alpha_{0}/\alpha}t^{\alpha_{0}/\alpha}\left\{1+R^{*}_{i}(t)+O(t^{-1/\alpha})\right\}\\ q_{j}(t)&\sim\tau c^{\prime}_{j}{}^{\alpha_{0}/\alpha}t^{\alpha_{0}/\alpha}\left\{1+R^{*}_{j}(t)+O(t^{-1/\alpha})\right\}\end{split}$ (A.13)
    as $t\rightarrow\infty$, in which $R^{*}_{i}(t)$ and $R^{*}_{j}(t)$ have the forms given in Corollary A.1.1. Consequently, we obtain: